Solve recurring relation :an+2 -10an+1+25an=5n(n≥0)

1
Expert's answer
2021-07-28T17:15:11-0400

Let us solve the recurrence relation "a_{n+2} -10a_{n+1}+25a_n=5^n,\ (n≥0)." The characteristic equation "k^2-10k+25=0" of homogeneous relation is equivalent to "(k-5)^2=0," and hence it has the roots "k_1=k_2=5." Therefore, the solution of the non-homogeneous recurrence relation is "a_n=(C_1+C_2n)5^n+p_n," where the particular solution of non-homogeneous recurrence relation is "p_n=an^25^n."


It follows that

"a(n+2)^25^{n+2}-10a(n+1)^25^{n+1}+25an^25^n=5^n," which is equivalent to

"25a(n^2+4n+4)5^{n}-50a(n^2+2n+1)5^{n}+25an^25^n=5^n," and hence

"25an^2+100an+100a-50an^2-100an-50a+25an^2=1."

We conclude that "50a=1," and thus "a=\frac{1}{50}."


Consequently, the solution of the recurrence relation "a_{n+2} -10a_{n+1}+25a_n=5^n" is

"a_n=(C_1+C_2n)5^n+\frac{1}{50}n^25^n" or "a_n=(C_1+C_2n)5^n+\frac{1}{2}n^25^{n-2}."